a body is thrown vertically upwards with initial velocity of 100m/s find (1)time taken to reach at maximum height (2)maximum height attend (3)velocity after 5 second (4)velocity after 16 second. Take (a=10m/s2)
1
Expert's answer
2012-07-24T11:25:02-0400
(1) time taken to reach at maximum height
tm = V0 / g = 100 / 10 = 10
s
(2) maximum height attend
h = V0 * t - g * tm^2 / 2 = 100 * 10 - 10 *
10^2 / 2 = 1000 - 1000 / 2 = 500 m
(3) velocity after 5 second
t = 5 <
tm = 10, so:
V(5) = V0 - g * t = 100 - 10 * 5 = 50 m/s
(4) velocity after
16 second
t = 16 > tm = 10, so:
V(16) = 0 + g * (t - tm) = 0 + 10 * (16
- 10) = 60 m/s